OpenStudy (anonymous):

How do I find all Zeros and write in polynominal factored form for: f(x)=x^4-256

OpenStudy (campbell_st):

this is \[f(x) = (x^2)^2 - 16^2\] so the problem is the difference of 2 squares so it can be factorised as \[(x^2 + 16)(x^2 - 16) = (x^2 +16)(x -4)(x+4)\] so the zeros are x = 4, -4 and \[x = \pm \sqrt{-16} ... or x = \pm 4i\]

OpenStudy (anonymous):

thank you...

OpenStudy (campbell_st):

hope it helps